Hassock Corp. produces woven wall hangings. It takes 2 hours of direct labor to produce a single wall hanging. Hassock’s standard labor cost is $12 per hour. During August, Hassock produced 19,000 units and used 38,140 hours of direct labor at a total cost of $454,680. What is Hassock’s labor rate variance for August?
Answer #1

Labor rate variance

= (SR - AR) *AH

= (12 - 454,680/38,410) * 38,140

= (12 - 11.92)*38,140

= 3000 FAVOURABLE

I have rounded some numbers. Comment if you face any issues